Connect Homework - Chapter 11 Saved Help Save & Exit Submit Retained earnings you, v restricted by treasury STOCK Less cost o I need help with the Explanation and Solution for the whole question Please and Thank you.
Answer #1

2) Cash dividend

Jan 5 Apr 5 July 5 Oct 5
Cash dividend 35000*.50 = 17500 (35000-5000)*.50 = 15000 15000 (30000*1.2*.50) = 18000